OpenStudy (anonymous):

if x=2 ..what iz x ....x=2+2

OpenStudy (anonymous):

this question doesnt make sense. if x=2, then the second part says 2=2+2

OpenStudy (anonymous):

If x = 2, then x ^ x = 2 + 2, also x * x = 2 + 2, and x + x = 2 + 2

OpenStudy (anonymous):

nevermind ill try to figure it by myself